OpenStudy (anonymous):

write solution set and graph on a number line: lx+7l>or equal to 3

OpenStudy (anonymous):

I got x>or equal to -10 and -4. Is that right? and how do I graph that?

OpenStudy (anonymous):

\[\left| x+7 \right| \ge 3 \implies x+7 \le -3~~or~~x+7 \ge 3 \implies x \le -10~~or~~x \ge -4\]\[\implies x \in (-\infty,-10] \cup [-4, \infty)\]

OpenStudy (anonymous):

Ok I just forgot to flip the sign!

OpenStudy (anonymous):

For \[\left| A \right| \ge b\]we get a graph that looks likewhich gives us the inequalities with which we started this problem. If you draw the picture, it's easy to get the equations.